# Coupled Requirements-driven Testing of CPS in Simulation and Reality Coupled Requirements-driven Testing of CPS: From Simulation To Reality

### v1

@@ Abstract
Software Testing is a well-established research area, with mature processes, tools, and techniques [1]. Over the years, a multitude of testing methods, such as model-based testing [2], mutation testing [3], metamorphic testing [4], and search-based testing [5], to name but a few, have emerged. Furthermore, the ISO 29119 standard [6], for example, covers testing processes, documentation, techniques, and guidelines for thorough software application testing. Depending on the application domain and context, software testing contributes to detecting bugs and errors in the System under Test (SuT), uncovering performance issues, and identifying missing or incorrectly implemented requirements contributing to assessing the quality of a software product. In case of safety-critical applications, testing and test results are leveraged as evidence for safety assurance, arguing about system safety, for example, as part of Safety Assurance Cases (SACs) [7].
@@
Particularly for Cyber-Physical Systems (CPS), such as robotic applications, operating as part of smart production environments, or small Uncrewed Aerial Systems (sUAS), used in search-and-rescue operations, failures during operation can result in severe incidents harming physical infrastructure, the surrounding environment, or even humans. It is, therefore, crucial to specify safety requirements, implement mitigations, and moreover, establish a thorough testing and validation process [8]. In CPS, this is commonly achieved with a combination of traditional testing activities, simulations conducted in a high-fidelity simulator, as well as Hardware in the Loop (HITL), and field tests. While, for example, for the automotive domain, thorough test and safety standards have been established [9], other areas still lack a thorough testing and validation process. This is particularly true in the domain of sUAS and small unmanned ground vehicles (UGV). While a multitude of different simulation environments exist, providing support for testing individual aspects of a CPS, such as collision avoidance, path planning, and vision-based navigation, an overarching and holistic testing and validation process is still missing. Furthermore, in this context, environmental conditions, play a particular role, as these systems need to operate in real-world scenarios, with diverse combinations, of, for example, wind conditions, weather, lighting, terrain geographical information. Simulators, however, are typically limited in what can be simulated and tested, and the level of fidelity, of how well a simulation represents the real world and the behavior of the hardware, greatly varies depending on the simulation environment [10]. This results in situations where a system that has been thoroughly tested, with numerous simulation runs being executed, still deviates from its specified behavior when used in the “real world” due to insufficient or low-fidelity simulations being executed [11]. To this end, we have created an initial framework for testing CPS, with a particular emphasis on sUAS and robotic applications that strongly rely on simulations, HITL, and ultimately real-world field testing. Based on our research in the domain of sUAS applications, and the lack of stringent regulations in this area (e.g., compared to the automotive domain [9]), we found that there is a dearth of structured testing concepts that incorporate requirements, diverse validation properties, and multi-facet integration testing (e.g., different simulations, in conjunction with field tests within a structured well-defined test context). Based on these initial insights, and our own experiences, where we have faced similar issues, we have derived an initial testing process and framework.
@@
The contributions of this RE@Next paper are as follows: First, we discuss existing work and provide a motivating example for the need for additional research. Second, we lay out a structured testing process, and the artifacts/steps necessary in the different phases. Third, we provide an application example, of how the process would be executed for performing tests for an sUAS system. Fourth, we present a research roadmap of open questions and avenues that require additional research.
@@ II Background & Motivating Example
sUAS Simulation Tools and Capabilities: CPS in general, and sUAS in particular, are complex in nature, combining hardware and software components, and human interaction, under complex and uncertain environments. Wind, for example, is a significant environmental factor impacting sUAS operations, leading to a range of hazards and unpredictable behaviors [12]. For instance, Meta’s Aquila project, an initiative to develop a fixed-wing UAS capable of distributing internet, experienced a setback when the UAS could not adequately handle an unexpected wind gust, resulting in a severe crash [13]. The post-incident report and analysis indicated that the sudden wind gust changed the UAS’s angle of descent, causing its autopilot to malfunction.
@@
Despite the availability of a large number of specialized and open-source sUAS simulation tools, testing realistic windy conditions remains a challenge for developers. The nature of wind is highly dynamic, as its velocity and direction are influenced by the complex geometry of the environment. However, widely popular simulation tools, including Gazebo[14] and AirSim [15], only allow for the simulation of basic wind conditions. These include constant wind in a specific direction applying uniform wind pressure throughout the flight, or turbulent wind based on a statistical model without considering the complex geometry of the environment. These simplifications in the wind simulations create a significant gap between the testing results in simulations and the actual behavior of sUAS during live operations. Further, simulation tools overlook other crucial factors such as GPS signal fluctuations, communication quality degradation, and radio interference, which significantly affect the real-world sUAS operations.
@@
CPS-Testing: In the context of CPS, a wide variety of testing and validation approaches exist to validate the behavior of sUAS before field deployments. Techniques range from assessing the accuracy of onboard AI models, to applying software testing techniques such as Fuzzy and Metamorphic testing [4], to automatically generate new diverse test cases. However, most of these approaches have only focused on validating certain aspects of the SuT, often employed in isolation, focusing on individual subsystems or components, rather than the entire deployable sUAS. For instance, consider an autonomous sUAS designed for wildfire monitoring and surveillance. In such scenarios, a sUAS’s obstacle avoidance system must swiftly navigate around unpredictable obstacles in the air, while its imaging system must keep a steady focus on hotspots, despite the sUAS’ movements. Furthermore, environmental factors such as hard-to-predict wind gusts must be considered. Besides individual functionality, all components of the system need to operate flawlessly in unison, which is often left untested by standard testing procedures.
@@
In a recent survey, Abbaspour et al. [16] created a categorization and comparison of CPS testing levels and analyzed them in the context of the V-Model. Two of the main challenges that were uncovered are (1) “Extra Functional Properties” that need to be taken into account, and moreover (2) providing support for (automated) evaluation of the test results. The latter, however, is an integral aspect of a holistic testing process, particularly when simulation is involved, as typically a large number of simulation runs are executed, with hundreds or even thousands of configuration parameters involved. Particularly when testing in simulations, e.g., in a 3-D environment, it is not always immediately clear if and to what extent a simulation has in fact executed correctly, or if any (minor) deviations from the expected behavior occurred. Similar challenges for CPS testing were reported by Zhou et al. [17], further emphasizing uncertainties and the interactions with physical systems, and the need to “ simulate and co-simulate many parallel physical processes to analyze the continuous interaction of the user, controller, and physical environment and to automatically identify the worst-case scenarios”.
@@
Drawing from the challenges in existing CPS testing, and our experience with CPS and sUAS simulations, we have developed a preliminary framework. The next section outlines this framework, detailing the process, model, and potential applications.
@@ III-A Requirements Modelling
Additionally, pertaining to CPS in general, and sUAS applications in particular, their operation in close proximity to humans, in real-world environments requires several safety-related aspects, and hence safety requirements, to be taken into consideration. Therefore, to facilitate proper testing and validation, as a precursor, we assume the existence of a Requirements Model that documents these in a structured manner, and a Verification & Validation Model (V&V Model) that contains information on properties to be validated. In particular, these properties should include both (a) test properties such as maintaining minimum sUAS battery during the mission, keeping flight path deviation percentage under a threshold level, and not breaching a restricted airspace, and (b) environment properties such as windy conditions, lighting conditions, and GPS satellite availability in the environment under which the SuT should maintain those test properties. We do not prescribe a specific modeling technique or language, but any technique that facilitates property specification and validation criteria can be used. Examples include formal approaches such as temporal logic, UML-based notations [26], or state machines in combination with model-checking [27].
@@ III-B Test Design & Management
III-C Test Story Generation & Execution:
@@
Req Model |
Reqs. for the SuT. [formalized, structured, or NL] |
V&V Model |
Properties of the SuT that need to hold and should be tested. [formalized, e.g., temporal logic] |
Test Model |
Description of the scenarios that must be tested, LoF requirements. [state machines, structured text] |
Test Exec. Req. |
Required capabilites of the execution environment (e.g, required simulator features). [structured text, e.g., XML/JSON] |
Fixture |
Implementation and setup for SuT to execute tests. [code, setup scripts, parameterization files] |
Test Story |
"Materialization" of a concrete test to be executed (specific to the test context (e.g. simulator execution, or field test guide "handbook")[executable scripts and properties, guidelines] |
Test Trace |
Collected runtime testing data. [SuT log files, died test protocol] |
Test Report |
Aggregated and processed test results [structured and visualized results, automated test verdicts] |
Concrete test stories for the sUAS requirement, for example, specify the characteristics of the 3D environment such as distinguishing between urban and open settings, and defining the direction, type, and velocity of wind within that environment. Further, these scenarios detail the specific tasks or missions assigned to the sUAS, as well as the initial conditions including the home location of each sUAS in the environment.
@@ IV-A Testing sUAS Systems
R1: An sUAS shall complete a flight with multiple waypoints in wind gusts
R2:An sUAS shall complete a flight without colliding with stationary objects, the terrain, or other aircraft
@@
R1 |
wind_gusts <= 23mph & deviation <5% |
FinalState: State active ‘‘prearm-checks successful’’ GoToState ready-for-takeoff State ready-for-takeoff ‘‘mission-assigned’’ GoToState request-takeoff ... |
Geospatial Data |
Wind-model |
(Vel, Dir, Coord) |
Simulator: Airsim |
Sim Env: Wind Config |
SuT sUAS Mission : JSON |
SuT-Sim Conn.: String |
Run-time Monitors: Scripts |
Flight Path |
sUAS IMU States |
SuT States |
Resultant Forces on sUAS |
Claim: sUAS does not collide with objects in dense environment when flying under 23mph gusty winds SC1: sUAS remains stable under gusty winds of up to 23 mph. SC2: sUAS avoids collision in dense environment (buildings, ground, other sUAS). |
R2 |
obs_density=0.4 & col.count =0 & miss.success = True |
Geospatial data |
Wind-model |
(Vel, Dir, Coord) |
Obstacle |
(Type, Location) |
Simulator: Airsim |
Sim Env: Cesium (Unreal) |
SuT sUAS Mission : JSON |
SuT-Sim Conn.: String |
Run-time Monitors: Scripts |
Landing Location |
sUAS IMU States |
SuT States |
Proximity to Obstacles |
@@
Following the successful execution of test story at LoF-2, developers can finally proceed to execute the same test story in the field (LoF-3). During field test, developers need to ensure that the testing site complies with the environmental requirements such as wind velocity in the area and obstacle density, and accordingly plan the SuT sUAS flight path. Developers should be able to reuse the same monitoring scripts used in LoF-2 to verify the test properties. Furthermore, developers should also collect hardware logs, such as temperature readings, power supply voltage, GPS strength, and radio signal reception, to provide additional context to the test artifacts. This process of gradually moving from a lower to a higher fidelity testing environment enables the comparison of test results between levels and provides feedback to the lower fidelity testing environment for refinements.
@@ IV-B Discussion
Our preliminary experimental application has demonstrated that our process is flexible enough to be applied to different testing activities, covering the need for both simulation and field testing activities. Based on these initial insights, a structured test model, along with clearly specified test execution requirements provides a formal framework to determine which tests need to be executed in which context and what properties or conditions need to be considered. Further, a clearly specified test "handbook" (compiled from the test model), with steps and actions explicitly outlined, will lead to documented, reproducible test results.
Generalizability: Concerning the application of our process beyond sUAS systems, further evaluations need to be conducted. However, based on our experience with robotic applications, we are confident that the process is also applicable to systems with similar characteristics. For example, testing an autonomous robot in a smart warehouse with requirements such as "When a human worker enters the path of the robot, the Computer Vision (CV) component shall detect the human and avoid a collision." could follow a similar testing process. The V&V model in this instance contains properties related to the detection speed and accuracy, as well as avoidance states. For execution requirements, the simulation environment must include a 3D model of a worker. The model must reflect a wide range of human characteristics, such as age, sex, height, skin tone, and clothing preferences, to ensure diversity. Additionally, the model must be programmed to mimic the typical behaviors of a warehouse worker [34]. Such execution requirements, for example, result in Gazebo [14] being selected as simulation environment for certain test stories in conjunction with CV tests. Using a ROS-based application, used for robotic applications, monitors can easily be established via topic subscriptions, allowing to collect and log runtime data and analyze the results of the simulations as evidence for an SAC about safe navigation in the presence of a human worker.
Future Evaluation Activities: We are actively pursuing two directions. First, we plan on applying an extended and iteratively refined process (cf. research roadmap) to a specific application scenario from our sUAS system (e.g., a search-and-rescue or surveillance scenario) where we fully track the testing process end to end, and subsequently apply it to a second system in the robotics domain, as sketched out above. Secondly, we plan to conduct a user study with sUAS developers and testers to evaluate the process through simulation and field testing, gaining insights into the effectiveness of various formalisms for models and artifacts.

### v2

@@ Abstract
Software Testing is a well-established research area, with mature processes, tools, and techniques [1]. Over the years, a multitude of testing methods, such as model-based testing [2], mutation testing [3], metamorphic testing [4, 5], and search-based testing [6, 7], to name but a few, have emerged. Furthermore, the ISO 29119 standard [8], for example, covers testing processes, documentation, techniques, and guidelines for thorough software application testing. Depending on the application domain and context, software testing contributes to detecting bugs and errors in the System under Test (SuT), uncovering performance issues, and identifying missing or incorrectly implemented requirements contributing to assessing the quality of a software product. In case of safety-critical applications, testing and test results are leveraged as evidence for safety assurance, arguing about system safety, for example, as part of Safety Assurance Cases (SACs) [9, 10].
@@
Particularly for Cyber-Physical Systems (CPS), such as robotic applications, operating as part of smart production environments, or small Uncrewed Aerial Systems (sUAS), used in search-and-rescue operations, failures during operation can result in severe incidents harming physical infrastructure, the surrounding environment, or even humans. It is, therefore, crucial to specify safety requirements, implement mitigations, and moreover, establish a thorough testing and validation process [11]. In CPS, this is commonly achieved with a combination of traditional testing activities, simulations conducted in a high-fidelity simulator, as well as Hardware in the Loop (HITL), and field tests. While, for example, for the automotive domain, thorough test and safety standards have been established [12], other areas still lack a thorough testing and validation process. This is particularly true in the domain of sUAS and small unmanned ground vehicles (UGV). While a multitude of different simulation environments exist, providing support for testing individual aspects of a CPS, such as collision avoidance, path planning, and vision-based navigation, an overarching and holistic testing and validation process is still missing. Particularly due to the fact that thorough validation and/or certification is not mandated by standards or regulations, these types of systems often only employ ad-hoc testing strategies. Furthermore, in this context, environmental conditions, play a particular role, as these systems need to operate in real-world scenarios, with diverse combinations, of, for example, wind conditions, weather, lighting, terrain geographical information. Simulators, however, are typically limited in what can be simulated and tested, and the level of fidelity, of how well a simulation represents the real world and the behavior of the hardware, greatly varies depending on the simulation environment [13]. This results in situations where a system that has been thoroughly tested, with numerous simulation runs being executed, still deviates from its specified behavior when used in the “real world” due to insufficient or low-fidelity simulations being executed [14].
To this end, we have created an initial framework for testing CPS, with a particular emphasis on sUAS and robotic applications that strongly rely on simulations, HITL, and ultimately real-world field testing. Based on our research in the domain of sUAS applications, and the lack of stringent regulations in this area (e.g., compared to the automotive domain [12]), we found that there is a dearth of structured testing concepts that incorporate requirements, diverse validation properties, and multi-facet integration testing (e.g., different simulations, in conjunction with field tests within a structured well-defined test context). Based on these initial insights, and our own experiences, where we have faced similar issues, we have derived an initial testing process and framework.
@@
The contributions of this RE@Next paper are as follows: First, we discuss existing work and provide a motivating example for the need for additional research. Second, we lay out a structured testing process, and the artifacts/steps necessary in the different phases. Third, we provide an application example, of how the process would be executed for performing tests for a sUAS system. Fourth, we present a research roadmap of open questions and avenues that require additional research.
@@ II Background & Motivating Example
sUAS Simulation Tools and Capabilities: CPS in general, and sUAS in particular, combine hardware, software, and human interaction components, under complex and uncertain environments. Wind, for example, is a significant environmental factor impacting sUAS operations, leading to a range of hazards and unpredictable behaviors [15]. For instance, Meta’s Aquila project, an initiative to develop a fixed-wing UAS capable of distributing internet, experienced a setback when the UAS could not adequately handle an unexpected wind gust, resulting in a severe crash [16].
@@
Despite numerous specialized and open-source sUAS simulation tools, realistically testing sUAS behavior in simulated windy conditions is challenging. The nature of wind is highly dynamic, as its velocity and direction are influenced by the complex geometry of the environment. However, widely popular simulation tools, including Gazebo[17] and AirSim [18], only allow for the simulation of basic wind conditions. These include constant wind in a specific direction applying uniform wind pressure throughout the flight, or turbulent wind based on a statistical model without considering the complex geometry of the environment. These simplifications in the wind simulations create a significant gap between the testing results in simulations and the actual behavior of sUAS during live operations. Further, simulation tools overlook other crucial factors such as GPS signal fluctuations, communication quality degradation, and radio interference, which significantly affect the real-world sUAS operations.
@@
CPS-Testing: In the context of CPS, a wide variety of testing and validation approaches exist to validate the behavior of sUAS before field deployments. Techniques range from assessing the accuracy of onboard AI models, to applying software testing techniques such as Fuzzy and Metamorphic testing [4], to automatically generate new diverse test cases. However, most of these approaches have only focused on validating certain aspects of the SuT, often employed in isolation, focusing on individual subsystems or components, rather than the entire deployable sUAS. For instance, consider an autonomous sUAS designed for wildfire monitoring and surveillance. In such scenarios, a sUAS’s obstacle avoidance system must swiftly navigate around unpredictable obstacles in the air, while its imaging system must keep a steady focus on hotspots, despite the sUAS’ movements. Besides individual functionality, all components of the system need to operate flawlessly in unison, which is often left untested by standard testing procedures.
@@
In a recent survey, Abbaspour et al. [19] created a categorization and comparison of CPS testing levels and analyzed them in the context of the V-Model. Two of the main challenges that were uncovered are (1) “Extra Functional Properties” that need to be taken into account, and moreover (2) providing support for (automated) evaluation of the test results. The latter, however, is an integral aspect of a holistic testing process, particularly when simulation is involved, as typically a large number of simulation runs are executed, with hundreds or even thousands of configuration parameters involved. For instance, during simulation testing, it is not always immediately clear if, and to what extent, a simulation has in fact executed correctly, or if any (minor) deviations from the expected behavior occurred. Similar challenges for CPS testing were reported by Zhou et al. [20], further emphasizing uncertainties and the interactions with physical systems, and the need to “ simulate and co-simulate many parallel physical processes to analyze the continuous interaction of the user, controller, and physical environment and to automatically identify the worst-case scenarios”.
@@
The goal of our work is to leverage these existing techniques to advance and streamline the testing process. Drawing from the challenges in existing CPS testing, and our experience with CPS and sUAS simulations, we have developed a preliminary framework. The next section outlines this framework, detailing the process, model, and potential applications.
@@ III-A Requirements Modelling
Additionally, pertaining to CPS in general, and sUAS applications in particular, their operation in close proximity to humans, in real-world environments, requires several safety-related aspects, and hence safety requirements, to be taken into consideration. Therefore, to facilitate proper testing and validation, as a precursor, we assume the existence of a Requirements Model that documents these in a structured manner, and a Verification & Validation Model (V&V Model) that contains information on properties to be validated. In particular, these properties should include both (a) test properties such as maintaining minimum sUAS battery during the mission, keeping flight path deviation percentage under a threshold level, and not breaching a restricted airspace, and (b) environment properties such as windy conditions, lighting conditions, and GPS satellite availability in the environment under which the SuT should maintain those test properties. We do not prescribe a specific modeling technique or language, but any technique that facilitates property specification and validation criteria can be used. Examples include formal approaches such as temporal logic, UML-based notations [29], or state machines in combination with model-checking [30].
@@ III-B Test Design & Management
III-C Test Story Generation & Execution
@@
Req Model | Reqs. for the SuT. [formalized, structured, or NL] |
V&V Model | Properties of the SuT that need to hold and should be tested. [formalized, e.g., temporal logic] |
Test Model | Scenario descriptions that must be tested, LoF requirements. [state machines, structured text] |
Test Exec. Req. | Required capabilities of the execution environment (e.g., required simulator features). [structured text, e.g., XML/JSON] |
Fixture | Implementation and setup for SuT to execute tests. [code, setup scripts, parameterization files] |
Test Story | "Materialization" of a concrete test to be executed specific to the test context (e.g., simulator execution, or field test guide "handbook").[executable scripts and properties, guidelines] |
Test Trace | Collected runtime testing data. [SuT log files, died test protocol] |
Test Report | Aggregated and processed test results. [structured and visualized results, automated test verdicts] |
Concrete test stories for the sUAS requirement, for example, specify the characteristics of the 3D environment, such as distinguishing between urban and open settings, and defining the direction, type, and velocity of wind within that environment. Further, these scenarios detail the specific tasks or missions assigned to the sUAS, as well as the initial conditions including the home location of each sUAS in the environment.
@@ IV-A Testing sUAS Systems
R1: A sUAS shall complete a flight with multiple waypoints in wind gusts
R2:A sUAS shall complete a flight without colliding with stationary objects, the terrain, or other aircraft
@@
R1 | wind_gusts <= 23mph & deviation <5% | FinalState: State active ‘‘prearm-checks successful’’ GoToState ready-for-takeoff State ready-for-takeoff ‘‘mission-assigned’’ GoToState request-takeoff ... | Geospatial Data | |||
Wind model | ||||||
(Vel, Dir, Coord) | ||||||
Simulator: AirSim | ||||||
Sim Env: Wind Config | ||||||
SuT sUAS Mission: JSON | ||||||
SuT-Sim Conn.: String | ||||||
Runtime Monitors: Scripts | ||||||
Flight Path | ||||||
sUAS IMU States | ||||||
SuT States | ||||||
Resultant Forces on sUAS | ||||||
Claim: sUAS does not collide with objects in dense environment when flying under 23mph gusty winds
SC1: sUAS remains stable under gusty winds of up to 23 mph SC2: sUAS avoids collision in dense environment (buildings, ground, other sUAS) |
R2 | obs_density=0.4 & col.count =0 & miss.success = True | Geospatial data | ||||
Wind model | ||||||
(Vel, Dir, Coord) | ||||||
Obstacle | ||||||
(Type, Location) | Simulator: AirSim | |||||
Sim Env: Cesium (Unreal) | ||||||
SuT sUAS Mission: JSON | ||||||
SuT-Sim Conn.: String | ||||||
Runtime Monitors: Scripts | Landing Location | |||||
sUAS IMU States | ||||||
SuT States | ||||||
Proximity to Obstacles | ||||||
@@
Following the successful execution of a test story at LoF-2, developers can finally proceed to execute the same test story in the field (LoF-3). During field tests, developers need to ensure that the testing site complies with the environmental requirements such as wind velocity in the area and obstacle density, and accordingly plan the SuT sUAS flight path. Developers should be able to reuse the same monitoring scripts used in LoF-2 to verify the test properties. Furthermore, developers should also collect hardware logs, such as temperature readings, power supply voltage, GPS strength, and radio signal reception, to provide additional context to the test artifacts. This process of gradually moving from a lower to a higher fidelity testing environment enables the comparison of test results between levels and provides feedback to the lower fidelity testing environment for refinements.
@@ IV-B Discussion
Our preliminary experimental application has demonstrated that our process is adaptable to various testing activities, and suitable for both simulation and field testing. Based on these initial insights, a structured test model with clear test execution requirements provides a formal framework to identify tests, execution context, and monitoring properties. Further, a detailed test "handbook" derived from the model, outlining steps and actions, results in documented, reproducible tests.
Generalizability: Concerning the application of our process beyond sUAS systems, further evaluations need to be conducted. However, based on our experience with robotic applications, we are confident that the process is also applicable to systems with similar characteristics. For example, testing an autonomous robot in a smart warehouse with requirements such as "When a human worker enters the path of the robot, the Computer Vision (CV) component shall detect the human and avoid a collision" could follow a similar testing process. The V&V model in this instance contains properties related to the detection speed, accuracy, and avoidance states. The simulation environment must include a 3D model of a worker that reflects diverse human characteristics – age, sex, height, skin tone, and clothing preferences – and mimics typical warehouse worker behaviors to meet execution requirements. Such execution requirements, for example, result in Gazebo [17] being selected as simulation environment for certain test stories in conjunction with CV tests. Using a ROS-based application, used for robotic applications, a monitoring system can easily be established to collect and log runtime data and analyze the results of the simulations as evidence for an SAC about safe navigation in the presence of a human worker.
Future Evaluation Plans: We are actively pursuing two directions. First, we plan on applying an extended and iteratively refined process (cf. research roadmap) to a specific application scenario from our sUAS system (e.g., a search-and-rescue or surveillance scenario) where we fully track the testing process end to end, and subsequently apply it to a second system in the robotics domain, as sketched out above. Secondly, we plan to conduct a user study with sUAS developers to evaluate this process through both simulation and field testing, aiming to assess the effectiveness of various modeling formalisms.